OpenStudy (anonymous):

Please help! If f^-1(x) exists, find this function algebraically. f(x)=-x^3 f^-1(x)

OpenStudy (anonymous):

rewrite to y=-x^3 to find the inverse, you'd replace x with y, y with x, and solve for y x=-y^3 y^3=-x y=(-x)^(1/3)=f^-1

OpenStudy (anonymous):

Hmmm.. Here are my options though :/ a. does not exist b. -x^3 c. =-1/x^3 d. -3\[\sqrt{x}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

It keeps saying you're typing but nothing is showning up :/

OpenStudy (anonymous):

*showing

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

don't know what to tell you, f^-1(x)=-x^(1/3) is the inverse of f(x)=-x^3 you can show this by composition f(f^-1(x))=f^-1(f(x))=x

OpenStudy (anonymous):

O.k. Thanks for your help Dockworker :)

OpenStudy (anonymous):

f(x)=-x^3 and is a one-to-one function, so it's inverse is definitely a function

OpenStudy (anonymous):

yw